OpenStudy (anonymous):

the perimeter of a rectangle is 60 feet. how do you find its length and width if the length is 8 feet longer than the width?

OpenStudy (nikato):



OpenStudy (nikato):

remember perimeter=sum of all sides

OpenStudy (nikato):

so w+w+8+w+8+w=60

OpenStudy (nikato):

now solve for w

OpenStudy (anonymous):

4w+8=60

OpenStudy (anonymous):

w=11

OpenStudy (anonymous):

and length=19

OpenStudy (anonymous):

2(19)+2(11)=60 feet

OpenStudy (nikato):

correct!

OpenStudy (anonymous):

thank you again (:

OpenStudy (nikato):

no problem. happy to help